OpenStudy (anonymous):

If f(x) = -2x + 1, find f(-6). 12 13 -11 -7

OpenStudy (studygurl14):

f(x) = -2x + 1 f(-6) = -2(-6) + 1 f(-6) = 12 + 1 f(-6) = 13